A rectangular plot has the length of 6 meters longer than it's width. The area of the plot is 16 square meters. Find the length and width of the plot.

Answer by jorel555(1290)

Let w be width. Then length is w+6. So:
w(w+6)=16
w²+6w-16=0
(w+8)(w-2)=0
w=2
w+6=8
The plot is 2m by 8m. ☺☺☺☺